OpenStudy (anonymous):

The weekly salaries (in dollars) for 8 employees of a small business are given below. (Note that these are already ordered from least to greatest.) 588 , 749 , 751 , 784 , 796 , 820 , 856 , 936 . Suppose that the $588 salary changes to $748. Answer the following. (A) what happens to the median? -it decrease by by $ -it increase by $ -it stays the same. (B) what happens to the mean? -it decrease by by $ -it increase by $ -it stays the same.

OpenStudy (kira_yamato):

748 , 749 , 751 , 784 , 796 , 820 , 856 , 936 Median remains the same because position did not change. Mean increased.

OpenStudy (anonymous):

What is the number of mean?

OpenStudy (kira_yamato):

Sample mean = ∑x/n = (748 + 749 + 751 + 784 + 796 + 820 + 856 + 936)/8 = 805

OpenStudy (texaschic101):

I agree...median stays the same and mean increases